# Coarse-to-Fine Latent Diffusion for Pose-Guided Person Image Synthesis

### 28 Feb 2024 v1

@@ Abstract
Diffusion model is a promising approach to image generation and has been employed for Pose-Guided Person Image Synthesis (PGPIS) with competitive performance. While existing methods simply align the person appearance to the target pose, they are prone to overfitting due to the lack of a high-level semantic understanding on the source person image. In this paper, we propose a novel Coarse-to-Fine Latent Diffusion (CFLD) method for PGPIS. In the absence of image-caption pairs and textual prompts, we develop a novel training paradigm purely based on images to control the generation process of the pre-trained text-to-image diffusion model. A perception-refined decoder is designed to progressively refine a set of learnable queries and extract semantic understanding of person images as a coarse-grained prompt. This allows for the decoupling of fine-grained appearance and pose information controls at different stages, and thus circumventing the potential overfitting problem. To generate more realistic texture details, a hybrid-granularity attention module is proposed to encode multi-scale fine-grained appearance features as bias terms to augment the coarse-grained prompt. Both quantitative and qualitative experimental results on the DeepFashion benchmark demonstrate the superiority of our method over the state of the arts for PGPIS. Code is available at https://github.com/YanzuoLu/CFLD.
@@ 1 Introduction
Figure 1: (a) The appearance of person image varies significantly given only a textual prompt for image generation by using Stable Diffusion [32] or ControlNet [47] with OpenPose guidance [3]. (b) Simply aligning the source appearance to the target pose without a semantic understanding of person image can easily lead to overfitting, such that the generated images become distorted and unnatural. (c) Our method learns the coarse-grained prompt for a comprehensive perception of the source image and injects fine-grained appearance features as bias terms, thus generating high-quality images with better generalization performance.
@@ 2 Related Work
2.1 Pose-Guided Person Image Synthesis
Ma et al. [23] first presents the task of pose-guided person image synthesis and refines the generated images in an adversarial manner. To decouple the pose and appearance information, early approaches [6, 24] propose to learn pose-irrelevant features but fail to handle the complex texture details with vanilla convolutional neural networks. To alleviate this problem, auxiliary information is introduced to improve the generation quality, such as parsing [25] and UV maps [35]. Recent approaches [54, 39, 17, 30, 19, 31] focus on modeling the spatial correspondence between pose and appearance, with the more frequent use of parsing maps [46, 22, 53]. PIDM [1] and PoCoLD [9] are developed based on diffusion models to prevent from the drawbacks in the generative adversarial networks, including the instability of min-max training objective and difficulty in synthesising high-resolution images. Both of these two diffusion-based methods extend the idea of spatial correspondence to model the relation between the appearance of source image and target pose via the cross-attention mechanism. We argue this leads to overfitting by simply aligning the source appearance to the target pose without a high-level semantic understanding of the person image. In this work, we propose a coarse-to-fine latent diffusion method by incorporating both the coarse-grained prompt and fine-grained appearance features. The coarse-grained prompt is gradually refined to give a semantic perception, while the fine-grained appearance features are injected as bias terms to control texture details in our proposed hybrid-granularity attention module. Both the quantitative and qualitative experiments validate that our method achieves state-of-the-art results with stronger generalization performance.
@@ 2.1 Pose-Guided Person Image Synthesis
2.2 Controllable Diffusion Models
Diffusion models have recently emerged and demonstrated their potential for high-resolution image synthesis. The core idea is to start with a simple noise vector and gradually transform it into a high-quality image through multiple denoising iterations. Beyond unconditional generation [15, 37, 38], various methods have been introduced to incorporate user-supplied control signals into the generation process, enabling more controllable image generation. For instance, [5] introduces the usage of classifier gradients to condition on the generation, while [14] proposes a classifier-free control mechanism employing a weighted summation of conditional and unconditional outputs for controllable synthesis. Moreover, the Latent Diffusion Model (LDM) performs diffusion in the latent space and injects the conditioning signals via a specific encoder and cross-attention. Building upon the pre-trained LDM like Stable Diffusion (SD) [32], subsequent works have explored to bias the latent space by adding extra controls [47, 26], as well as further to provide users with control over the generated content [40, 12]. Rather than employing a high-level conditioning prompt throughout the generation, we design a coarse-to-fine conditioning process that adjusts the latent features at different stages within the UNet-based prediction network, providing better controllable pose-guided person image synthesis.
@@ 3.3 Optimization
When the source image $\bm{x}_{s}$ is missing, we use learnable vectors as the conditional embeddings. The learnable vectors are trained with a probability of $\eta$to drop both $\bm{x}_{s}$ and $\bm{x}_{p}$ during training. The outputs of the pose adapter $\mathcal{H}_{P}$ will be set to all zeros if the target pose $\bm{x}_{tp}$ is missing. We use the DDIM scheduler [37] to speed up the sampling with 50 steps as the same as in [9, 1]. Finally, the generated image is obtained by the VAE decoder $\bm{y}=\mathcal{D}(\tilde{\bm{z}}_{0})$.
@@ 4 Experiments
Table 2: Quantitative comparisons with the state of the arts in terms of image quality. ${}^{\dagger}$ We strictly follow the evaluation implementation in NTED [31] and reproduce these results based on the checkpoints rather than generated images provided by the authors for a fair comparison. ${}^{\ddagger}$ Results are obtained using the generated images released by the authors. $*{}^{*}$ Results are cited from PoCoLD [9] without publicly available checkpoints and generated images. Others are cited from NTED.
Figure 3: Qualitative comparisons with SPGNet [22], DPTN [48], NTED [31], CASD [53] and PIDM [1].
@@ 4.3 Qualitative Comparison
In Fig. 3, we present a comprehensive visual comparison with recent approaches that are publicly available and reproducible [^1] The checkpoints and generated images of the related work PoCoLD [9] are not publicly available for qualitative comparison., including SPGNet [22], DPTN [48], NTED [31], CASD [53] and PIDM [1]. Our observations can be summarized as follows. (1) Both GAN-based and diffusion-based methods suffer from overfitting the human poses. When generating some target poses that are extreme or not common in the training set, existing methods show severe distortions as demonstrated in rows 1-2. Since we decouples the controls of fine-grained appearance and pose information, our method circumvents the potential overfitting problem and always generates a reasonable pose with the conditioning coarse-grained prompt and fine-grained appearance bias. (2) For source images in rows 3-6 with more complex clothing, our generated images better preserve the textures details while aligning with the target pose thanks to the robust coarse-to-fine learning curriculum of hybrid-granularity attention module. For other methods, although they match in color, the clothes either exhibit blurring and distortion (SPGNet, DPTN, and CASD) or are spliced unnaturally in texture, creating a large gap from the source image (NTED and PIDM). (3) As for cases where the target pose requires visualization of areas invisible in the source image, our method exhibits strong understanding and generalization capabilities. With a semantic understanding of the source image provided by the perception-refined decoder, our method is aware of what should be predicted when the person turns around or sits down as illustrated in rows 7-10, such as different patterns on the front and back of clothes, the sitting chair, and lower body wear.
@@ 5 Conclusion
This paper presents a novel Coarse-to-Fine Latent Diffusion (CFLD) method for Pose-Guided Person Image Synthesis (PGPIS). We circumvent the potential overfitting problem by decoupling the fine-grained appearance and pose information controls. Our proposed Perception-Refined Decoder (PRD) and Hybrid-Granularity Attention module (HGA) enable a high-level semantic understanding of person images, while also preserving texture details through a coarse-to-fine learning curriculum. Extensive experiments demonstrate that CFLD outperforms the state of the arts in PGPIS both quantitatively and qualitatively. Our future work will investigate whether the CFLD can be extended to more general image synthesis tasks besides PGPIS.

### 09 Apr 2024 v2

@@ Abstract
Diffusion model is a promising approach to image generation and has been employed for Pose-Guided Person Image Synthesis (PGPIS) with competitive performance. While existing methods simply align the person appearance to the target pose, they are prone to overfitting due to the lack of a high-level semantic understanding on the source person image. In this paper, we propose a novel Coarse-to-Fine Latent Diffusion (CFLD) method for PGPIS. In the absence of image-caption pairs and textual prompts, we develop a novel training paradigm purely based on images to control the generation process of a pre-trained text-to-image diffusion model. A perception-refined decoder is designed to progressively refine a set of learnable queries and extract semantic understanding of person images as a coarse-grained prompt. This allows for the decoupling of fine-grained appearance and pose information controls at different stages, and thus circumventing the potential overfitting problem. To generate more realistic texture details, a hybrid-granularity attention module is proposed to encode multi-scale fine-grained appearance features as bias terms to augment the coarse-grained prompt. Both quantitative and qualitative experimental results on the DeepFashion benchmark demonstrate the superiority of our method over the state of the arts for PGPIS. Code is available at https://github.com/YanzuoLu/CFLD.
@@ 1 Introduction
Figure 1: (a) The appearance of person images varies significantly given only a textual prompt for image generation by using Stable Diffusion [35] or ControlNet [55] with OpenPose guidance [3]. (b) Simply aligning the source appearance to the target pose without a semantic understanding of person image can easily lead to overfitting, such that the generated images become distorted and unnatural. (c) Our method learns the coarse-grained prompt for a comprehensive perception of the source image and injects fine-grained appearance features as bias terms, thus generating high-quality images with better generalization performance.
@@ 2 Related Work
Pose-Guided Person Image Synthesis. Ma et al. [26] first presents the task of pose-guided person image synthesis and refines the generated images in an adversarial manner. To decouple the pose and appearance information, early approaches [6, 27] propose to learn pose-irrelevant features but fail to handle the complex texture details with vanilla convolutional neural networks. To alleviate this problem, auxiliary information is introduced to improve the generation quality, such as parsing [28] and UV maps [38]. Recent approaches [63, 44, 18, 33, 20, 34] focus on modeling the spatial correspondence between pose and appearance, with the more frequent use of parsing maps [54, 25, 62]. PIDM [1] and PoCoLD [9] are developed based on diffusion models to prevent from the drawbacks in the generative adversarial networks, including the instability of min-max training objective and difficulty in synthesising high-resolution images. Both of these two diffusion-based methods extend the idea of spatial correspondence to model the relation between the appearance of source image and target pose via the cross-attention mechanism. We argue this leads to overfitting by simply aligning the source appearance to the target pose without a high-level semantic understanding of the person image. More concurrent works like MagicAnimate [50], Animate Anyone [16] and PCDMs [39] require multi-stage and progressive fully fine-tuning, while our pipeline is more efficient and end-to-end by freezing most parameters. And the training paradigm of IP-Adatper [52] is heavily relying on image-text pairs which are not available for PGPIS task.
@@
Controllable Diffusion Models. Diffusion models have recently emerged and demonstrated their potential for high-resolution image synthesis. The core idea is to start with a simple noise vector and gradually transform it into a high-quality image through multiple denoising iterations. Beyond unconditional generation [15, 42, 43], various methods have been introduced to incorporate user-supplied control signals into the generation process, enabling more controllable image generation. For instance, [5] introduces the usage of classifier gradients to condition on the generation, while [14] proposes a classifier-free control mechanism employing a weighted summation of conditional and unconditional outputs for controllable synthesis. Moreover, the Latent Diffusion Model (LDM) performs diffusion in the latent space and injects the conditioning signals via a specific encoder and cross-attention. Building upon the pre-trained LDM like Stable Diffusion (SD) [35], subsequent works have explored to bias the latent space by adding extra controls [55, 29], as well as further to provide users with control over the generated content [45, 12]. Rather than employing a high-level conditioning prompt throughout the generation, we design a coarse-to-fine conditioning process that adjusts the latent features at different stages within the UNet-based prediction network, providing better controllable pose-guided person image synthesis.
@@ 3.3 Optimization
When the source image $\bm{x}_{s}$ is missing, we use learnable vectors as the conditional embeddings. The learnable vectors are trained with a probability of $\eta$to drop both $\bm{x}_{s}$ and $\bm{x}_{p}$ during training. The outputs of the pose adapter $\mathcal{H}_{P}$ will be set to all zeros if the target pose $\bm{x}_{tp}$ is missing. We use the DDPM scheduler [15] with 50 steps as the same as in [9, 1]. Finally, the generated image is obtained by the VAE decoder $\bm{y}=\mathcal{D}(\tilde{\bm{z}}_{0})$.
@@ 4 Experiments
Table 2: Quantitative comparisons with the state of the arts in terms of image quality. ${}^{\dagger}$ We strictly follow the evaluation implementation in NTED [34] and reproduce these results. ${}^{\ddagger}$ Results are obtained using the generated images released by the authors.
Figure 3: Qualitative comparisons with state-of-the-arts. To clarify the relation between objective and subjective metrics, we demonstrate the LPIPS measures and label the images with the first and second highest votes from user opinions as red and blue respectively.
@@ 4.3 Qualitative Comparison
In Fig. 3, we present a comprehensive visual comparison with recent approaches that are publicly available and reproducible, including SPGNet [25], DPTN [56], NTED [34], CASD [62] and PIDM [1]. Our observations can be summarized as follows. (1) Both GAN-based and diffusion-based methods suffer from overfitting the human poses. When generating some target poses that are extreme or not common in the training set, existing methods show severe distortions as demonstrated in rows 1-2. Since we decouples the controls of fine-grained appearance and pose information, our method circumvents the potential overfitting problem and always generates a reasonable pose with the conditioning coarse-grained prompt and fine-grained appearance bias. (2) For source images in rows 3-6 with more complex clothing, our generated images better preserve the textures details while aligning with the target pose thanks to the robust coarse-to-fine learning curriculum of hybrid-granularity attention module. For other methods, although they match in color, the clothes either exhibit blurring and distortion (SPGNet, DPTN, and CASD) or are spliced unnaturally in texture, creating a large gap from the source image (NTED and PIDM). (3) As for cases where the target pose requires visualization of areas invisible in the source image, our method exhibits strong understanding and generalization capabilities. With a semantic understanding of the source image provided by the perception-refined decoder, our method is aware of what should be predicted when the person turns around or sits down as illustrated in rows 7-10, such as different patterns on the front and back of clothes, the sitting chair, and lower body wear.
@@ 5 Conclusion
This paper presents a novel Coarse-to-Fine Latent Diffusion (CFLD) method for Pose-Guided Person Image Synthesis (PGPIS). We circumvent the potential overfitting problem by decoupling the fine-grained appearance and pose information controls. Our proposed Perception-Refined Decoder (PRD) and Hybrid-Granularity Attention module (HGA) enable a high-level semantic understanding of person images, while also preserving texture details through a coarse-to-fine learning curriculum. Extensive experiments demonstrate that CFLD outperforms the state of the arts in PGPIS both quantitatively and qualitatively. Our future work will investigate whether the CFLD can be extended to more downstream tasks that suffer from inferior data like person re-identification [23, 53] and domain adaptation [24, 40], since our training paradigm yields both a pre-trained feature network and a powerful generator for augmentation.
Acknowledgments. This work was supported in part by the National Natural Science Foundation of China (U22A2095, 62276281).